# IdealTSF: Can Non-Ideal Data Contribute to Enhancing the Performance of Time Series Forecasting Models?

### v1

@@ Introduction
Time series forecasting tasks are ubiquitous across various domains, including economics (Granger and Newbold 2014), energy (Martín et al. 2010; Qian et al. 2019), transportation planning (Chen et al. 2001; Yin et al. 2021), weather forecasting (Wu et al. 2023), healthcare, and natural sciences. Over the past decade, deep learning methods (LeCun, Bengio, and Hinton 2015) have gained popularity in forecasting, often outperforming statistical methods such as ARIMA (Shumway and Stoffer 2017). However, until recently, deep learning methods for forecasting have primarily focused on selecting superior feature extraction techniques, utilizing specific training schemes, and employing foundational architectures such as CNNs (Wang et al. 2023; Wu et al. 2022; Hewage et al. 2020), RNNs (Lai et al. 2018; Qin et al. 2017), Transformers (Vaswani et al. 2017; Zhou et al. 2022; Nie et al. 2022), and MLP variants (Challu et al. 2023; Murad, Aktukmak, and Yilmaz 2025), with the aim of capturing the trend characteristics of time series data as effectively as possible.
@@
Conventional time series analysis methods typically operate under the assumption that data are complete and devoid of anomalies. However, this assumption frequently fails to hold true in real-world applications (Schmidl, Wenig, and Papenbrock 2022; Qu et al. 2024). Common approaches to handling missing values, such as interpolation, typically presume a linear or smooth relationship between missing entries and observed data (Chen et al. 2023). Such methods tend to oversimplify the inherent complexity of time series data, neglecting potential nonlinear dependencies, abrupt fluctuations, and various non-stationary dynamics (Xu et al. 2022; Chen et al. 2024). For instance, in meteorological datasets, certain weather patterns may shift as a consequence of climate change, rendering simplistic interpolation techniques inadequate for capturing such evolving trends. In the presence of large-scale missing data, the efficacy of interpolation techniques becomes markedly limited, often failing to reconstruct the underlying data distribution accurately (Zeng et al. 2025a, b). Conversely, conventional outlier detection techniques frequently depend on statistical assumptions, such as normality of the data distribution (Zhang et al. 2025b, a; Yao, Li, and Xiao 2024). Such assumptions may lead to the misclassification of meaningful fluctuations as anomalies, thereby compromising the predictive performance of the model. For example, in power load forecasting, seasonal patterns or exceptional events—such as national holidays—can induce substantial fluctuations in load data (Xiao et al. 2024; Wang et al. 2025). Traditional anomaly detection methods may erroneously label these fluctuations as outliers, thereby overlooking critical trend information and introducing bias into forecasts (Xiao et al. 2025; Yao et al. 2023; Lu et al. 2025).
@@
In practical applications, time series data often exhibit unpredictable and non-ideal variations (Zhang et al. 2025c, 2024). Simply compensating for data imperfections may impair the model’s generalization ability. Therefore, enhancing the model’s robustness to imperfect conditions is critical. To this end, we propose IdealTSF, a time series forecasting framework designed to jointly strengthen the model’s resilience across the pre-training, training, and optimization phases. As illustrated in Figure 2, the data exhibit pronounced jumps and heavy-tailed characteristics. IdealTSF explicitly addresses these non-ideal properties by introducing a negative sample pre-training module, which utilizes stable distributions to generate probabilistic data with targeted statistical features—mimicking jump processes and heavy-tailed stochastic behaviors. Additionally, it incorporates multi-scale noise injection and structured deletion to simulate anomalies and missingness commonly observed in real-world events. During training, we generate positive samples using hybrid smoothed interpolation, and integrate them with original inputs through a pre-trained attention mechanism to extract predictive features. Finally, in the optimization stage, negative perturbations are introduced to guide gradient descent toward flatter minima. To further enhance generalization, adversarial training is employed using FGSM (Fast Gradient Sign Method) or PGD (Projected Gradient Descent) attacks. This enables the model to converge more rapidly to flat optimal solutions and remain resilient to imperfect data. As demonstrated in Figure 1, IdealTSF outperforms state-of-the-art deep models, achieving approximately a 10% improvement in optimization metrics.
@@ Methods
In this section, we present the experimental results of IdealTSF on several mainstream long-term series forecasting (LTSF) benchmark datasets, covering both long- and short-term forecasting tasks for multivariate time series. To further validate the effectiveness of negative samples in IdealTSF, we conduct additional experiments that involve pre-training solely on negative samples without using any positive samples.
@@
Table 1 presents a comparison between IdealTSF and other baseline models across eleven benchmark datasets. The results show that IdealTSF consistently ranks among the top two on all datasets, often achieving or closely approaching state-of-the-art performance. These datasets span various series with different sampling frequencies, numbers of variables, and real-world application scenarios. Notably, IdealTSF significantly outperforms models like Twinformer (Zhou et al. 2025), TimeMixer++ (Wang et al. 2024), TimeKAN (Huang et al. 2025), SAMformer (Ilbert et al. 2024), and PatchTST (Nie et al. 2022), and, relative to TimeKAN, reduces MSE on ECL and ETTh1 by approximately 17% and 3.5%, respectively. Importantly, IdealTSF also demonstrates robust performance on datasets with inherently low predictability, such as Traffic and ECL, further validating its generalization ability. In addition to long-term forecasting, IdealTSF achieves strong results in short-term forecasting tasks as well. On the PEMS benchmark—which consists of multiple time series recorded across a city-wide traffic network—many advanced models (e.g., PatchTST (2023) and DLinear (2023)) exhibit performance degradation due to complex spatiotemporal dependencies among variables. In contrast, IdealTSF maintains competitive performance on this challenging task, highlighting its effectiveness in modeling complex multivariate time series.
@@
Figure 7: The left figure shows the performance of the Adam and ECOS+Adam optimizers. The right figure presents the performance of the SGD and ECOS+SGD optimizers.

### v2

@@ Introduction
Time series forecasting tasks are ubiquitous across various domains, including economics (granger2014forecasting), energy (martin2010prediction; qian2019review), transportation planning (chen2001freeway; yin2021deep), weather forecasting (wu2023interpretable), healthcare, and natural sciences. Over the past decade, deep learning methods (lecun2015deep) have gained popularity in forecasting, often outperforming statistical methods such as ARIMA (shumway2017arima). However, until recently, deep learning methods for forecasting have primarily focused on selecting superior feature extraction techniques, utilizing specific training schemes, and employing foundational architectures such as CNNs (wang2023micn; wu2022timesnet; hewage2020temporal), RNNs (lai2018modeling; qin2017dual), Transformers (vaswani2017attention; zhou2022fedformer; nie2022time), and MLP variants (challu2023nhits; murad2025wpmixer), with the aim of capturing the trend characteristics of time series data as effectively as possible.
@@
Conventional time series analysis methods typically operate under the assumption that data are complete and devoid of anomalies. However, this assumption frequently fails to hold true in real-world applications (schmidl2022anomaly; qu2024livescene). Common approaches to handling missing values, such as interpolation, typically presume a linear or smooth relationship between missing entries and observed data (chen2023imdiffusion). Such methods tend to oversimplify the inherent complexity of time series data, neglecting potential nonlinear dependencies, abrupt fluctuations, and various non-stationary dynamics (xu2022tgan; chen2024freegaussian). For instance, in meteorological datasets, certain weather patterns may shift as a consequence of climate change, rendering simplistic interpolation techniques inadequate for capturing such evolving trends. In the presence of large-scale missing data, the efficacy of interpolation techniques becomes markedly limited, often failing to reconstruct the underlying data distribution accurately (zeng2025FSDrive; zeng2025janusvln). Conversely, conventional outlier detection techniques frequently depend on statistical assumptions, such as normality of the data distribution (zhang2025thatsn; zhang2025probabilistic; yao2024swift). Such assumptions may lead to the misclassification of meaningful fluctuations as anomalies, thereby compromising the predictive performance of the model. For example, in power load forecasting, seasonal patterns or exceptional events—such as national holidays—can induce substantial fluctuations in load data (xiao2024confusion; wang2025c3). Traditional anomaly detection methods may erroneously label these fluctuations as outliers, thereby overlooking critical trend information and introducing bias into forecasts (xiao2025diffusion; yao2023ndc; lu2025mcnr).
@@
In practical applications, time series data often exhibit unpredictable and non-ideal variations (zhang2025enhancing; zhang2024cf). Simply compensating for data imperfections may impair the model’s generalization ability. Therefore, enhancing the model’s robustness to imperfect conditions is critical. To this end, we propose IdealTSF, a time series forecasting framework designed to jointly strengthen the model’s resilience across the pre-training, training, and optimization phases. As illustrated in Figure 2, the data exhibit pronounced jumps and heavy-tailed characteristics. IdealTSF explicitly addresses these non-ideal properties by introducing a negative sample pre-training module, which utilizes stable distributions to generate probabilistic data with targeted statistical features—mimicking jump processes and heavy-tailed stochastic behaviors. Additionally, it incorporates multi-scale noise injection and structured deletion to simulate anomalies and missingness commonly observed in real-world events. During training, we generate positive samples using hybrid smoothed interpolation, and integrate them with original inputs through a pre-trained attention mechanism to extract predictive features. Finally, in the optimization stage, negative perturbations are introduced to guide gradient descent toward flatter minima. To further enhance generalization, adversarial training is employed using FGSM (Fast Gradient Sign Method) or PGD (Projected Gradient Descent) attacks. This enables the model to converge more rapidly to flat optimal solutions and remain resilient to imperfect data. As demonstrated in Figure 1, IdealTSF outperforms state-of-the-art deep models, achieving approximately a 10% improvement in optimization metrics.
@@ Methods
In this section, we present the experiments of IdealTSF on multiple mainstream datasets and the effectiveness tests of relevant modules. To prove the robustness of IdealTSF against non - ideal real - world factors, we conducted experiments on all baseline models in short - term prediction by randomly adding noise ranging from 5% to 35% again. Meanwhile, to ensure fairness, we presented the complete experimental data without noise addition in the appendix.
@@
Table 1 presents a comparison between IdealTSF and other baseline models across eleven benchmark datasets. The results show that IdealTSF consistently ranks among the top two on all datasets, often achieving or closely approaching state-of-the-art performance. These datasets span various series with different sampling frequencies, numbers of variables, and real-world application scenarios. Notably, IdealTSF significantly outperforms models like Twinformer (zhoutwinsformer), TimeMixer++ (wang2024timemixer++), TimeKAN (huang2025timekan), SAMformer (ilbert2024samformer), and PatchTST (nie2022time), and, relative to TimeKAN, reduces MSE on ECL and ETTh1 by approximately 17% and 3.5%, respectively. Importantly, IdealTSF also demonstrates robust performance on datasets with inherently low predictability, such as Traffic and ECL, further validating its generalization ability. In addition to long-term forecasting, IdealTSF achieves strong results in short-term forecasting tasks as well. On the PEMS benchmark—which consists of multiple time series recorded across a city-wide traffic network—many advanced models (e.g., PatchTST (2023) and DLinear (2023)) exhibit performance degradation due to complex spatiotemporal dependencies among variables. In contrast, IdealTSF maintains competitive performance on this challenging task, highlighting its effectiveness in modeling complex multivariate time series.
@@
Figure 7: (a) The left figure shows the performance of the Adam and ECOS+Adam optimizers. (b) The right figure presents the performance of the SGD and ECOS+SGD optimizers.